# Quantum automata and languages of finite index

### v2

@@ Abstract
This paper continues the study of measure-once finite quantum automata building on work by Bertoni et al. and Blondel et al. We investigate conditions ensuring that, given a language recognized by such a device and a language generated by a context-free grammar of finite index or by a matrix context-free grammar, it is decidable whether or not they have a nonempty intersection.
Keywords: Quantum automata, Context-free languages, Algebraic groups, Decidability
AMS subject classifications: 81P68, 68Q45
@@ 1 Introduction
In this paper, we investigate an issue of decidability for quantum (finite) automata. Quantum automata were introduced at the beginning of 2000’s in [23] as a new model of language recognizer (cf. [5]). Numerous publications have ever since compared their decision properties to those of other models and, in particular, to probabilistic finite automata. In this theoretical setting, we investigate the decidability of a problem for quantum automata introduced in [6, 7], named Intersection Problem. The starting point of that investigation is a result of Blondel, Jeandel, Koiran, and Portier who showed in [8] that, given a (“measure-once” in the sense of [23]) quantum automaton $\cal Q$, it is decidable whether the language recognized by $\cal Q$ has a non trivial intersection with the free monoid generated by the input alphabet of $\cal Q$ (Emptiness problem). It is worth remarking that these problems are strictly related to some computational issues concerning matrices. Precisely, a crucial step in the construction of the decision procedure for both the Emptiness and Intersection Problem amounts to effectively compute the Zariski closure of a finitely generated group of matrices over $\mathbb{Q}$. For that task, an appropriate algorithm has been developed in [12] by Derksen, Jeandel, and Koiran.
Before coming back to the Intersection Problem, we find then convenient to recall some developments of the research on that aspect. Two striking results have been obtained recently in connection with the algorithm mentioned above. In 2022 Nosan, Pouly, Schmitz, Shirmohammadi, and Worrell, following an alternative approach for such computation, have obtained in [25] a bound on the degree of the polynomials that define the Zariski closure of the set and showed that the latter can be computed in elementary time.
The second was obtained in 2023 by Hrushovski, Ouaknine, Pouly, and Worrell for the study of the effective computation of polynomial invariants for affine programs [17]. Precisely, it is exhibited an algorithm that, given a finite set of square matrices of the same size with coefficients in $\mathbb{Q}$, computes the Zariski closure of the semigroup generated by the set.
@@
where $s\in\mathbb{R}^{n}$ is a row-vector of unit Euclidean norm, $P$ is a projection matrix of $\mathbb{R}^{n}$, $\varphi:\Sigma^{*}\rightarrow O_{n}$ is a morphism from the free monoid $\Sigma^{*}$, generated by the input alphabet $\Sigma$ of $\cal Q$, into the group $O_{n}$ of orthogonal $n\times n$-matrices over $\mathbb{R}$ and the threshold $\lambda$ has value in $\mathbb{R}$.
We are mainly interested in effective properties which require the quantum automaton to be effectively given. We thus consider the model of rational quantum automaton, i.e., where all the coefficients of the components of (1) and $\lambda$ are rational numbers. For a real threshold $\lambda$, the languages recognized by $\cal Q$ with strict and nonstrict threshold $\lambda$ are
$|{\cal Q}_{>}|=\{w\in\Sigma^{*}\mid||s\varphi(w)P||>\lambda\},\quad|{\cal Q}_{\geq}|=\{w\in\Sigma^{*}\mid||s\varphi(w)P||\geq\lambda\},$
@@
As already pointed out, the decidability of the above-mentioned problem has been proven in [8] in the specific case $L$ $=\Sigma^{*}$. This result contrasts with the corresponding problems (strict and non strict, respectively) for probabilistic finite automata. Indeed for this class of automata, the Emptiness problems w.r.t. $|{\cal Q}_{>}|$ and $|{\cal Q}_{\geq}|$ are both undecidable (see [27], Thm 6.17). Also it contrasts with the undecidability of the Emptiness problem of $|{\cal Q}_{\geq}|$ for quantum automata [8]. It is also proven in [21] that both the problems (strict and non strict, respectively) remain undecidable for the “measure many” model of quantum automata, a model not computationally equivalent to the “measure once”, introduced by Kondacs and Watrous in [22].
In order to describe the main contributions of the paper, it may be useful to recall some lines of the technique used in the proof of the Intersection Problem in [7]. For this purpose, following the same strategy of [8], we first observe that the Intersection Problem is equivalent to the inclusion
@@
$M\rightarrow||sMP||,$
– where $M$ is an arbitrary matrix in $\mathbb{R}^{n\times n}$, and $s,P$ are components of (1) – is continuous, it is sufficient to prove that, for all matrices $M$ in the Euclidean closure ${\mathbf{C}l}(\varphi(L))$ of $\varphi(L)$, the condition $||sMP||\leq\lambda$ holds (see Section 3). Such condition is then effectively tested by running, in parallel, two semialgorithms. The first tests the inclusion $\varphi(L)\subseteq|{\cal Q}_{>}|$ and it is readily seen to be semidecidable. The second, which is the main crux, is to prove that (2) is semidecidable.
The semidecidability of the latter procedure is achieved by expressing the property (2) in first-order logic of the field of reals, which, in turn, consists to effectively compute a representation of ${\mathbf{C}l}(\varphi(L))$ in term of a semialgebraic set (a more general property than algebraic, which allows one to cope with the fact that algebraic sets are not closed w.r.t. the operation of product of sets). Afterwards, one applies on the constructed formula the Tarski-Seidenberg quantifier elimination to verify whether (2) holds true or not.
@@
The reduction process mentioned above is made possible by resorting to two ingredients: the fact that the Euclidean closure ${\mathbf{C}l}(M_{A})$ is an algebraic set (because of a deep result of Algebraic Geometry, namely Theorem 3) and because of a suitably defined effective structuring of the computations, i.e., the derivations, of $G$ (Sections 4.1, 4.2 and 6.1). In this theoretical setting, the work in [7] and this work provide the following picture of the Intersection Problem:
@@
  • If, for every variable $A$ of $G$, $M_{A}$ is finitely generated, then ${\mathbf{C}l}(\varphi(L))$ is effectively computable. This is guaranteed by applying Proposition 9 and the results [12, 17, 25] mentioned above. This, in particular, covers the cases of context-free linear languages, bounded semi-linear languages and a new class of languages, proposed in this paper, generated by restricted matrix context-free linear grammars which non trivially extends the previous ones (Proposition 10).
  • The effective computation of ${\mathbf{C}l}(\varphi(L))$ is achieved for a subfamily of finite index context-free languages, called monoidal, by considering the property of (algebraic) irreducibility on the Zariski closure of a family of monoids of cycles canonically associated with $G$ (Proposition 11).
The new results of the paper are the following. In Section 6.1, we extend the main contribution of [7] – i.e., the statement of decidability of the Intersection Problem for linear context-free languages and for bounded semilinear languages – to a class of matrix context-free grammar of finite index. Precisely, whenever the quantum automaton is rational, such result (Proposition 10) allows one, on one hand, to recover both the previous two cases as immediate corollaries of such result. On the other, it shows the decidability of the problem for languages of exponential growth that are not context-free. In the proof a suitable decomposition of the derivations in these grammars is used. Such a decomposition is based upon some combinatorial tools proposed by Choffrut, Varricchio and the second author in [10] for the study of rational relations. In particular, this decomposition permits a refinement of the machinery developed in [7] for the manipulation of the monoids of cycles of these grammars.
In Section 6.2, we investigate the Intersection Problem for languages defined by finite index context-free grammars. These are grammars $G$ where each word $w$ in the language generated by $G$ is obtained by a derivation $\delta$, whose index is uniformly bounded, i.e., the number of variables in each sentential form of $\delta$ is bounded by an integer not depending on $w$. A remarkable result by Ginsburg and Spanier [14] (see also Nivat [24]) provides a charaterization of such languages in terms of composition of grammars: precisely, each language of this type is generated by a grammar $G$ given by the composition
@@
We prove that, regardless of the length $k$ of the composition (3) for $G$, if the Zariski closure of the monoids of cycles of the linear grammars of the lowest level ${\cal G}_{k}$ of (3), are (algebraic) irreducible, then one can effectively compute the Euclidean closure ${\mathbf{C}l}(\varphi(L))$ of $\varphi(L)$ (Proposition 11, Corollary 5).
Extensions of such result to broader families of finite index context-free languages seem related to possible extensions of the algorithms [12, 17, 25] (cf Section 7, [$\alpha$]). The general case of context-free languages is still open and it would be even possible that the problem is undecidable (cf Section 7, [$\beta$]).
The paper is structured in the following way. In Sec. 2 preliminaries on semialgebraic and irreducible algebraic sets are presented. In Section 3 we recall an effective condition to which the decidability of the problem under discussion can be reduced. Section 4 provides combinatorial structurings of context-free languages in the general case and in the finite index case. Section 5 recalls some results relating context-free grammars and semialgebraic sets. In Section 6 the main results of the paper are presented. Section 7 contains some concluding remarks and open problems. Section 8 shows the construction for monoidal grammars on an example.
@@ 2 Preliminaries on semialgebraic sets
A quantum automaton ${\cal Q}$ is a quadruple $(s,\varphi,P,\lambda)$ where $s\in\mathbb{R}^{n}$ is a vector of unit norm, $P$ is a projection of $\mathbb{R}^{n}$, i.e., $P\in\{0,1\}^{n\times n}$ with $P^{2}=P$, and $\varphi$ is a morphism
@@ 2.1 Topology
Consequently, if ${\cal A}$ is a binary relation which is a direct product ${\cal A}_{1}\times{\cal A}_{2}$, we have ${\mathbf{C}l}({\cal A}_{1}{\cal A}_{2})=f({\mathbf{C}l}({\cal A}_{1}\times{\cal A}_{2}))$. It is an elementary result of topology that ${\mathbf{C}l}({\cal A}_{1}\times{\cal A}_{2})={\mathbf{C}l}({\cal A}_{1})\times{\mathbf{C}l}({\cal A}_{2})$ holds. Because of ${\mathbf{C}l}({\cal A}_{1}{\cal A}_{2})=f({\mathbf{C}l}({\cal A}_{1}\times{\cal A}_{2}))=f({\mathbf{C}l}({\cal A}_{1})\times{\mathbf{C}l}({\cal A}_{2}))={\mathbf{C}l}({\cal A}_{1})\ {\mathbf{C}l}({\cal A}_{2})$ we have
@@ 2.2 Algebraic and semialgebraic sets
A subset ${\cal A}\subseteq\mathbb{R}^{n}$ is algebraic (over the field of real numbers), if ${\cal A}$ is the zero set of an arbitrary subset $\cal P$ of polynomials of $\mathbb{R}[x_{1},\ldots,x_{n}]$, i.e., for every vector $v\in\mathbb{R}^{n}$,
$v\in{\cal A}\ \Longleftrightarrow\ \forall\ p\in{\cal P}:\ p(v)=0.$ (5)
@@
$\sum_{p\in\cal P}p(x)^{2}=0.$
One can check that the family of algebraic sets is closed under finite unions and intersections. However, it is not closed under complement and projection. The following more general class of subsets enjoys extra closure properties and is therefore more robust.
@@
(ii) ${\cal A}$ is first-order definable in the theory of the structure whose domain are the reals and whose predicates are of the form $p(x_{1},\ldots,x_{n})>0$ and $p(x_{1},\ldots,x_{n})=0$ with $p\in\mathbb{R}[x_{1},\ldots,x_{n}]$.
@@
Notational convention. In connection with the $(L,{\cal Q})$ Intersection Problem, we will adopt the following terminology. A set ${\cal A}$ of matrices will be called effective algebraic (resp., effective semialgebraic) if the polynomials (resp., the formula) defining ${\cal A}$ can be algorithmically computed from the input of the problem, i.e., an effectively defined language $L$ and a finite quantum automaton $Q$.

### v3

@@ Abstract
In this paper, we provide some conditions of algebraic and combinatorial flavour that make decidable the following problem: given a measure-once finite quantum automaton and a context-free grammar of finite index or a simple matrix context-free grammar, the set intersection of the language accepted by the automaton and the language generated by the grammar, is empty or not.
Keywords: Quantum automata, Context-free languages, Algebraic groups, Matrix semigroups, Decidability
AMS subject classifications: 81P68, 68Q45, 03D05
@@ 1 Introduction
A fundamental issue in the mathematical theory of formal languages concerns the decidability of the Emptiness Problem: given a model of computation $\cal Q$, is the language accepted by $\cal Q$, $L({\cal Q})$, equal to the empty set? A natural generalization of such a problem, called Intersection Problem, arises by relativizing it in the following way: one considers a predefined family $\cal L$ of (effectively defined) languages and asks whether the intersection of $L({\cal Q})$ with a language taken arbitrarily from $\cal L$ is empty. Notice that, if $\cal L$ is the singleton set containing the free monoid of words over the input alphabet of $\cal Q$, one immediately obtains the Emptiness Problem.
In this paper, by continuing an investigation of Bertoni, Choffrut et al. [9, 10], we will study the decidability of the Intersection Problem w.r.t. quantum finite automata as the family $\cal Q$ and context-free languages, together with some other strictly related languages, as the family $\cal L$. Quantum automata were introduced at the beginning of 2000’s as a new model of language recognizer, that can be obtained by imposing the quantum paradigm on classical finite automata. Several models of quantum automata have been proposed in the literature, which essentially differ in what quantum measurements are allowed (see [1, 8, 21] for reference and updated texts). In this paper, we consider the model of a quantum automaton, called “measure-once”, introduced by Moore and Crutchfield in [33]. Such a model is the simplest and one of the most studied in the literature since it provides the description of good-featured quantum devices of “small size”. In this regard, it is worth recalling that a recent contribution of Mereghetti and Palano describes a method for the physical implementation of measure-once quantum automata for the recognition of a well-known family of periodic languages [31, 32].
Numerous publications have investigated the decision properties of quantum automata. The starting point of that investigation is a result of Blondel, Jeandel, Koiran, and Portier who showed in [11] that, given a quantum automaton $\cal Q$ (in the sense of [33]), it is decidable whether the language recognized by $\cal Q$ has a non trivial intersection with the free monoid generated by the input alphabet of $\cal Q$ (the Emptiness Problem, aforementioned). It is worth remarking that these problems are related to some computational issues concerning matrices. Precisely, a crucial step in the construction of the decision procedure for both the Emptiness and Intersection Problem amounts to effectively compute the Zariski closure of a finitely generated group of matrices over $\mathbb{Q}$. A striking result proved by Derksen, Jeandel and Koiran in [15] shows the existence of an appropriate algorithm for that task.
Before coming back to the Intersection Problem, we find then convenient to recall some developments of the research on that aspect. Two results have been obtained recently in connection with the algorithm of [15]. In 2022 Nosan, Pouly, Schmitz, Shirmohammadi, and Worrell, following an alternative approach for such computation, have obtained in [35] a bound on the degree of the polynomials that define the Zariski closure of the set and showed that the latter can be computed in elementary time.
The second was obtained in 2023 by Hrushovski, Ouaknine, Pouly, and Worrell for the study of the effective computation of polynomial invariants for affine programs [23, 22]. Precisely, it is exhibited an algorithm that, given a finite set of square matrices of the same size with coefficients in $\mathbb{Q}$, computes the Zariski closure of the semigroup generated by the set.
@@
where $s\in\mathbb{R}^{n}$ is a row-vector of unit Euclidean norm, $P$ is an orthogonal projection matrix of size ${n}$, $\varphi:\Sigma^{*}\rightarrow O_{n}$ is a morphism from the free monoid $\Sigma^{*}$, generated by the input alphabet $\Sigma$ of $\cal Q$, into the group $O_{n}$ of orthogonal $n\times n$-matrices over $\mathbb{R}$ and the threshold $\lambda$ has value in $\mathbb{R}$.
We are mainly interested in effective properties which require the quantum automaton to be effectively given. We thus consider the model of rational quantum automaton, i.e., where all the coefficients of the components of (1) and $\lambda$ are rational numbers. For a real threshold $\lambda$, the languages recognized by $\cal Q$ with strict and nonstrict threshold $\lambda$ are respectively
$|{\cal Q}_{>}|=\{w\in\Sigma^{*}\mid||s\varphi(w)P||^{2}>\lambda\},\quad|{\cal Q}_{\geq}|=\{w\in\Sigma^{*}\mid||s\varphi(w)P||^{2}\geq\lambda\},$
@@
As already pointed out, the decidability of this problem has been proven in [11] in the specific case $L$ $=\Sigma^{*}$. This result contrasts with the corresponding problems (strict and non strict, respectively) for probabilistic automata. Indeed, for these automata, the Emptiness Problems w.r.t. $|{\cal Q}_{>}|$ and $|{\cal Q}_{\geq}|$ are both undecidable [37], Thm 6.17 (see also [7] for the undecidability of the property to be isolated for a cut point). Also it contrasts with the undecidability of the Emptiness Problem of $|{\cal Q}_{\geq}|$ for quantum automata [11]. It is also proven in [29] that both the problems (strict and non strict, respectively) remain undecidable for the “measure many” model of quantum automata, a model not computationally equivalent to the “measure once”, introduced by Kondacs and Watrous in [30]. In order to describe the main contributions of the paper, it may be useful to recall some lines of the technique used in the proof of the Intersection Problem in [10]. For this purpose, following the same strategy of [11], we first observe that the Intersection Problem is equivalent to the inclusion
@@
$M\rightarrow||sMP||^{2},$
– where $M$ is an arbitrary matrix in $\mathbb{R}^{n\times n}$, and $s,P$ are components of (1) – is continuous, it is sufficient to prove that, for all matrices $M$ in the Euclidean closure ${\mathbf{C}l}(\varphi(L))$ of $\varphi(L)$, the condition $||sMP||^{2}\leq\lambda$ holds (see Section 3). Such condition is then effectively tested by expressing the property (2) in first-order logic of the field of reals, which, in turn, consists to effectively compute a representation of ${\mathbf{C}l}(\varphi(L))$ in term of a semialgebraic set (a more general property than algebraic, which allows one to cope with the fact that algebraic sets are not closed w.r.t. the operation of product of sets). Afterwards, one applies on the constructed formula the Tarski-Seidenberg quantifier elimination to verify whether (2) holds true or not.
@@
The reduction process mentioned above is made possible by resorting to two ingredients: the fact that the Euclidean closure ${\mathbf{C}l}(M_{A})$ is an algebraic set over the field $\mathbb{R}$ of real numbers, because of a deep result of Algebraic Geometry (see Theorem 3) and because of a suitably defined effective structuring of the computations, i.e., the derivations, of $G$ (Sections 4.1, 4.2 and 6.1). In this theoretical setting, the work in [10] and this work provide the following picture of the Intersection Problem:
@@
  • If, for every variable $A$ of $G$, $M_{A}$ is a regular submonoid, then ${\mathbf{C}l}(\varphi(L))$ is effectively computable. This is guaranteed by applying Proposition 9 and the results [15, 23, 35] mentioned above. This, in particular, covers the cases of context-free linear languages, bounded semi-linear languages and a new class of languages, proposed in this paper, generated by restricted matrix context-free grammars which non trivially extends the previous ones (Proposition 10).
  • The computation of ${\mathbf{C}l}(\varphi(L))$ is achieved for a subfamily of finite index context-free languages, called monoidal, by considering the property of (algebraic) irreducibility on the Zariski closure over $\mathbb{R}$ of a family of monoids of cycles canonically associated with $G$ (Proposition 11).
The new results of the paper are the following. In Section 6.1, we extend the main contribution of [10] to a subfamily of simple matrix languages. Simple matrix languages have been introduced by Ibarra in [24] as a proper class of the deterministic linear-bounded acceptor languages, defined by a restricted form of matrix context-free grammars (see [27] for recent developments). They contain, in particular, context-free languages and inherit several structural and decidable properties. Precisely, we prove that, whenever the quantum automaton is rational, such extension (Proposition 10) allows one, on one hand, to recover both the previous two cases – i.e., the case of linear context-free languages and that of bounded semi-linear languages – as immediate corollaries. On the other hand, it shows the decidability of the problem for languages of exponential growth that are not context-free. In the proof, a suitable decomposition of the derivations in these grammars is used. Such a decomposition is based upon some combinatorial tools proposed by Choffrut, Varricchio and the second author in [13] for the study of the counting functions of rational relations. In particular, this decomposition permits a refinement of the machinery developed in [10] for the manipulation of the monoids of cycles of these grammars.
In Section 6.2, we investigate the Intersection Problem for languages defined by finite index context-free grammars. These are grammars $G$ where each word $w$ in the language generated by $G$ is obtained by some derivation $\delta$, whose index is uniformly bounded, i.e., the number of variables in each sentential form of $\delta$ is bounded by an integer not depending on $w$. A result by Ginsburg and Spanier [18] (see also Nivat [34]) provides a characterization of such languages in terms of composition of grammars: precisely, each language of this type is generated by a grammar $G$ given by the composition
@@
We prove that, regardless of the length $k$ of the composition (3) for $G$, if the Zariski closure over $\mathbb{R}$ of the monoids of cycles of the grammars of the lowest level ${\cal G}_{k}$ of (3), are (algebraic) irreducible, then one can compute the Euclidean closure ${\mathbf{C}l}(\varphi(L))$ of $\varphi(L)$ (Proposition 11), thus making the Intersection Problem decidable (Corollaries 5 and 6).
Extensions of such result to broader families of finite index context-free languages seem related to possible extensions of the algorithms [15, 23, 35] (cf Section 7, [$\alpha$]). The general case of context-free languages is open and it could be even possible that the problem is undecidable (cf Section 7, [$\beta$]).
The paper is structured in the following way. In Section 2 preliminaries on semialgebraic and irreducible algebraic sets are presented. In Section 3 we recall an effective condition to which the decidability of the problem under discussion can be reduced. Section 4 provides combinatorial structurings of context-free languages in the general case and in the finite index case. Section 5 recalls some results relating context-free grammars and semialgebraic sets. In Section 6 the main results of the paper are presented. Section 7 contains some concluding remarks and open problems. In the Appendix, the construct of composition of grammars is given in detail.
The results of this paper have been announced at RP 2024 [5]. The present paper contains fully detailed proofs and examples of applications and construction.
@@ 2 Preliminaries on semialgebraic sets
A quantum automaton ${\cal Q}$ is a quadruple $(s,\varphi,P,\lambda)$ where $s\in\mathbb{R}^{n}$ is a vector of unit norm, $P$ is an orthogonal matrix of size $n$ with $P^{2}=P$, and $\varphi$ is a morphism
@@
Remark 1
Throughout the paper we will assume that the linear representation of every quantum automaton (with possibly complex entries) is given, as in Definition 1, by matrices over the field of real numbers. This is a merely technical restriction since any quantum automaton can be simulated by another one of the latter type. We refer the reader to the Section “Complex versus real entries” of the paper [11], where this issue is very clearly and thoroughly explained.
@@ 2.1 Topology
Thus, if ${\cal A}$ is a binary relation which is a direct product ${\cal A}_{1}\times{\cal A}_{2}$, we have ${\mathbf{C}l}({\cal A}_{1}{\cal A}_{2})=f({\mathbf{C}l}({\cal A}_{1}\times{\cal A}_{2}))$. It is an elementary result of topology that ${\mathbf{C}l}({\cal A}_{1}\times{\cal A}_{2})={\mathbf{C}l}({\cal A}_{1})\times{\mathbf{C}l}({\cal A}_{2})$ holds. Because of ${\mathbf{C}l}({\cal A}_{1}{\cal A}_{2})=f({\mathbf{C}l}({\cal A}_{1}\times{\cal A}_{2}))=f({\mathbf{C}l}({\cal A}_{1})\times{\mathbf{C}l}({\cal A}_{2}))={\mathbf{C}l}({\cal A}_{1})\ {\mathbf{C}l}({\cal A}_{2})$, we have the following
@@ 2.2 Algebraic and semialgebraic sets
A subset ${\cal A}\subseteq\mathbb{R}^{n}$ is algebraic (over the field of real numbers), if ${\cal A}$ is the zero set of an arbitrary subset $\cal P$ of polynomials of $\mathbb{R}[x_{1},\ldots,x_{n}]$, i.e., for every vector ${\bf v}\in\mathbb{R}^{n}$,
${\bf v}\in{\cal A}\ \Longleftrightarrow\ \forall\ p\in{\cal P}:\ p({\bf v})=0.$ (5)
@@
$\sum_{p\in\cal P}p({\bf v})^{2}=0.$
One can check that the family of algebraic sets is closed under finite unions and intersections. However, it is not closed under complement and projection. The following more general class of subsets enjoys extra closure properties and is thus more robust. The equivalence of the two definitions below is guaranteed by Tarski-Seidenberg quantifier elimination result (cf [4], Ch. 2, Sec. 2.3 and Sec. 2.5, Cor. 2.95).
@@
(ii) ${\cal A}$ is first-order definable in the theory of the structure whose domain are the reals and whose predicates are of the form $p(x_{1},\ldots,x_{n})>0$ or $p(x_{1},\ldots,x_{n})=0$ with $p\in\mathbb{R}[x_{1},\ldots,x_{n}]$.
@@
Notational convention. In connection with the $(L,{\cal Q})$ Intersection Problem, we will adopt the following terminology. A set ${\cal A}$ of matrices will be called effective algebraic (resp., effective semialgebraic) if the polynomials (resp., the formula) defining ${\cal A}$ can be algorithmically computed from the input of the problem, i.e., an effectively defined language $L$ and a finite quantum automaton $\cal Q$.